As a stockholder in Randolph Corporation, you receive its annual report. In the financial statements, Randolph has reported that the after-tax (net) income is $300 million. With 150 million shares of common stock outstanding, Randolph announced to distribute $100 million of dividends to its shareholders. The stock is now sold for $20 per share.
a. Assume that Randolph Corporation does not have any outstanding debt. The current share price reflects the fair value of the Corporation.
i. Find the market value of Randolph Corporation after the ex-dividend date,
ii. Find the price per share of Randolph Corporation after the ex-dividend date,
iii. Calculate the value of investor’s wealth who holds 4,000 shares of Randolph Corporation before the ex-dividend date,
1
Expert's answer
2019-12-05T09:40:36-0500

i. The market value of Randolph Corporation after the ex-dividend date is:

Dividend per share is: D = 100/150 = $0.66.

The price per share of Randolph Corporation after the ex-dividend date is:

20 - 0.66 = $19.33.

MV = 19.33*150,000,000 = $2.9 billion.

ii. The price per share of Randolph Corporation after the ex-dividend date is:

20 - 0.66 = $19.33.

iii. The value of investor’s wealth who holds 4,000 shares of Randolph Corporation before the ex-dividend date is:

$19.33*4,000 = $77,320.